OpenStudy (anonymous):

Please help me! Is this an inverse relation.... 2x+4y=6 Explain?

OpenStudy (anonymous):

We usually are asked if two functions are inverses, or find the inverse. Is there more to this question?

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

Have you been working with functions being inversely proportional?

OpenStudy (anonymous):

Um I don't think so.... we've been working with inverse variations.

OpenStudy (anonymous):

ok, that is what I needed. When a function varies inversely, it can be written as \[y = {k \over x}\] If we solve this function for y, we get \[2x+4y=6\]\[4y=-2x+6\]\[y = {-2 \over 4}x+{6 \over 4}\] which isn't in the same form as y = k/x so this function is not an example if inverse variation. OK?

OpenStudy (anonymous):

Oh thanks so much!